Substituting x = – 3 in
x^2 – 5x + 4 gives (– 3)^2 – 5( – 3)+4 = 9–15+4 = – 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Substituting x =–3 in x^2–5x+4
x^2–5x+4 = (–3)^2–5(–3)+4 = 9+15+4 = 28.
This is the correct answer

Thanks